# LLMs on the Line: Data Determines Loss-to-Loss Scaling Laws

### v1

@@ 6 Conclusion
We identify pretraining data and tokenizer as the dominant factors shaping these scaling laws, highlighting the importance of data curation. Architecture has limited impact, with models as different as LLaMA (transformer-based) and Mamba (a state-space model) exhibiting nearly identical scaling when trained on the same data and tokenizer. Model size, context length, and optimization settings have negligible influence, such that loss-to-loss scaling remains stable across different configurations.
@@ Acknowledgements
Appendix A Model Details
In addition to the models discussed in 4.1, we add additional details of Llama and Mamba models we trained of different depths and widths.
Table 1: Details of the models we trained from scratch using the tiktoken tokenizer.
@@ Appendix A Model Details
Appendix B Loss-to-Loss Scaling Across Settings
We supplement Fig. 2 from 3 with additional architecture-pretraining pairings in Figs. 10, 11, 12, 13, 14 and 15.
@@ Appendix B Loss-to-Loss Scaling Across Settings
Appendix C Intervention Results for Different Choices of x-Axis
We show variations of Figs. 4, 5 and 6 from 4.1 with C4 validation loss as the x-axis in Figs. 16, 17 and 18. Variations for Figs. 7, 8 and 9 from 4.2 are shown in Figs. 19, 20 and 21.
@@ Appendix C Intervention Results for Different Choices of x-Axis
Figure 17: The tokenizer has a moderate impact on loss-to-loss scaling laws.
@@
Appendix D Intervention Results without Averaging
The causal analysis in 4 was performed on scaling laws for average validation or test loss. Figs. 22, 23, 24, 25, 26 and 27 show illustrative results on scaling laws for individual datasets.
@@ Appendix D Intervention Results without Averaging
Figure 24: The tokenizer has a moderate impact on loss-to-loss scaling laws.
Figure 25: The tokenizer has a moderate impact on loss-to-loss scaling laws.
@@
Appendix E Additional Train-to-Test Scaling Laws
We provide train-to-test scaling laws for the interventions performed in 4.2 and Figs. 7, 8 and 9 in Figs. 31, 32, 33, 28, 29 and 30.
@@ Appendix E Additional Train-to-Test Scaling Laws
Figure 30: Optimizer settings do not affect train-to-test scaling.
Figure 31: Model size does not affect train-to-test scaling.

### v2

@@ 7 Conclusion
We identify pretraining data and as the dominant factor shaping these scaling laws, highlighting the importance of data curation. Architecture has limited impact, with models as different as LLaMA (transformer-based) and Mamba (a state-space model) exhibiting nearly identical scaling when trained on the same data and tokenizer. Tokenizer, model size, context length, and optimization settings have negligible influence, such that loss-to-loss scaling remains stable across different configurations.
@@ Acknowledgements
Appendix A Scaling Law Details
We adopt the compute-to-loss scaling law formulation from Brandfonbrener et al. (2024) Eq. (4):
$L\left(f^{(N,D)}\right)=E+\left(\left(\frac{A}{N}\right)^{\frac{\alpha}{\beta}}+\frac{B}{D}\right)^{\beta},$ (2)
where $f^{(N,D)}$ is a model with $N$ parameters trained on $D$ tokens and $E,A,B,\alpha,\beta$ are parameters to be fit. Notably, the irreducible error $E$ captures the minimum loss possible for model $f$ in the limit of infinite model and data size.
By default, Eq. 2 is fit using the training or validation loss. However, as demonstrated by Brandfonbrener et al. (2024) and our experiments, we can alternatively predict the loss $L_{x}$ on dataset $\mathcal{D}_{x}$ achieved by model $f_{p}^{(N,D)}$ trained on the pretraining set $\mathcal{D}_{p}$:
$L_{x}\left(f_{p}^{(N,D)}\right)=E_{x|p}+\left(\left(\frac{A}{N}\right)^{\frac{\alpha}{\beta}}+\frac{B}{D}\right)^{\beta}.$ (3)
As in Brandfonbrener et al. (2024) Eq. (7), the irreducible error
$E_{x|p}=L_{x}(f_{p}^{*})$ (4)
then captures the minimum possible loss on $\mathcal{D}_{x}$ of a model trained on $\mathcal{D}_{p}$.
With that, we can formulate the loss-to-loss scaling law for arbitrary combinations of pretraining data and two test or validation sets, as stated in Eq. 1.
Appendix B Fitting Details
As explained in 3, we fit loss-to-loss scaling law in Eq. 1 by collecting $200$ to $800$ model checkpoints throughout training and across model sizes and seeds.
For each line in a plot corresponding to a loss-to-loss scaling law from Eq. 1, we first fit the two compute-to-loss scaling laws $L_{x}(f_{p}^{(N,D)})$ and $L_{y}(f_{p}^{(N,D)})$ given by Eq. 3. This yields estimates for the irreducible errors $E_{x|p},E_{y|p}$, which correspond to the minimum x- and y-value of the loss-to-loss line. We use SciPy’s default curve_fit optimizer for fitting (Virtanen et al., 2020). In rare cases when all checkpoints have the same number of parameters $N$ or same number of tokens $D$ (this is the case only for a small subset of the Hugging Facemodels) and a compute-to-loss scaling law cannot be fitted, we instead estimate the irreducible error as the minimum loss achieved:
$E_{x|p}=\min_{N,D}L_{x}\left(f_{p}^{(N,D)}\right).$ (5)
We show example compute-to-loss fits for some of the loss-to-loss scaling laws from Fig. 2 in Fig. 10.
Figure 10: Example compute-to-loss scaling law fits. Each loss-to-loss scaling law requires fitting two compute-to-loss scaling laws to estimate $E_{x|p},E_{y|p}$. The three fits here are used for the The Pile UC and HellaSwag curves in Fig. 2, which showed curves for FineWeb-Edu-trained Mamba, i.e., $p$ is FineWeb-Edu. All curves in Fig. 2 use FineWeb-Edu as the x-axis; the corresponding $E_{x|p}$ is given by the fit depicted in the left-most plot. $E_{y|p}$ for $y$ as The Pile UC and HellaSwag are given by the fits in the center and right plot, respectively.
With $E_{x|p},E_{y|p}$ from the compute-to-loss fits, we again use SciPy’s curve_fit to fit $K,\kappa$ for the loss-to-loss scaling law from Eq. 1.
Appendix C Model Details
Table 2 supplements the discussion of the models used for our study from 4.1 with additional details.
Table 2: Architecture of the models we trained from scratch. All models generally use the tiktoken tokenizer. In addition to these models, for the tokenizer ablation 4.1, we train the largest Llama and Mamba models with the gpt2 tokenizer with beginning and end of sequence tokens.
@@ Appendix C Model Details
Table 3: Details of the pretrained models used. All models do not use beginning and end of sequence tokens.
Name Pretraining data Tokenizer Class Parameters
GPT-J The Pile GPT2 6B
GPT-Neo The Pile GPT2 {125M, 1.3B, 2.7B}
FineWeb Ablation Models The Pile GPT2 1.7B
FineWeb Ablation Models C4 GPT2 1.7B
FineWeb Ablation Models FineWeb-Edu GPT2 1.7B
GPT-NeoX The Pile GPT-NeoX 20B
Pythia (small) The Pile GPT-NeoX {70M, 160M, 410M}
Pythia (large) The Pile GPT-NeoX {1B, 1.4B, 2.8B, 6.9B, 12B}
Mamba The Pile GPT-NeoX {130M, 370M, 790M, 1.4B, 2.8B}
Mamba2 The Pile GPT-NeoX {130M, 370M, 790M, 1.4B, 2.8B}
Appendix D Quantitative Analysis of Interventions
We quantify the impact of different interventions as the area between fitted curves in Table 4. Pretraining data clearly has the biggest impact on the scaling laws.
Table 4: Area between fitted BPB curves for different interventions, evaluated on the interval [0, 2].
Base Tokenizer Data Architecture Optimizer Size Context
Base 0.01 0.22 0.02 0.02 0.02 0.02
Tokenizer 0.01 0.19 0.03 0.03 0.03 0.03
Data 0.22 0.19 0.17 0.19 0.19 0.19
Architecture 0.02 0.03 0.17 0.02 0.02 0.02
Optimizer 0.02 0.03 0.19 0.02 0.00 0.01
Size 0.02 0.03 0.19 0.02 0.00 0.00
Context 0.02 0.03 0.19 0.02 0.01 0.00
Appendix E Loss-to-Loss Scaling Across Settings
We supplement Fig. 2 from 3 with additional architecture-pretraining pairings in Figs. 11, 12, 13, 14, 15 and 16.
@@ Appendix E Loss-to-Loss Scaling Across Settings
Appendix F Intervention Results for Different Choices of x-Axis
We show variations of Figs. 4, 5 and 6 from 4.1 with C4 validation loss as the x-axis in Figs. 17, 18 and 19. Variations for Figs. 7, 8 and 9 from 4.2 are shown in Figs. 20, 21 and 22.
@@ Appendix F Intervention Results for Different Choices of x-Axis
Figure 18: The tokenizer has a minor impact on loss-to-loss scaling laws.
@@
Appendix G Intervention Results without Averaging
The causal analysis in 4 was performed on scaling laws for average validation or test loss. Figs. 23, 24, 25, 26, 27 and 28 show illustrative results on scaling laws for individual datasets.
@@ Appendix G Intervention Results without Averaging
Figure 25: The tokenizer has a minor impact on loss-to-loss scaling laws.
Figure 26: The tokenizer has a minor impact on loss-to-loss scaling laws.
@@
Appendix H Additional Train-to-Test Scaling Laws
We provide train-to-test scaling laws for the interventions performed in 4.2 and Figs. 7, 8 and 9 in Figs. 32, 33, 34, 29, 30 and 31.
Figure 29: Model size does not affect train-to-test scaling.
@@ Appendix H Additional Train-to-Test Scaling Laws
Figure 34: Optimizer settings do not affect train-to-test scaling.